### v2

# Worst-Case Symbolic Constraints Analysis and Generalisation with Large Language Models

Abstract.
Large language models (LLMs) have demonstrated strong performance on coding tasks such as generation, completion and repair, but their ability to handle complex symbolic reasoning over code still remains underexplored. We introduce the task of worst-case symbolic constraints analysis, which requires inferring the symbolic constraints that characterise worst-case program executions; these constraints can be solved to obtain inputs that expose performance bottlenecks or denial-of-service vulnerabilities in software systems. We show that even state-of-the-art LLMs (e.g., GPT-5) struggle when applied directly on this task. To address this challenge, we propose WARP, an innovative neurosymbolic approach that computes worst-case constraints on smaller concrete input sizes using existing program analysis tools, and then leverages LLMs to generalise these constraints to larger input sizes.
Concretely, WARP comprises: (1) an incremental strategy for LLM-based worst-case reasoning, (2) a solver-aligned neurosymbolic framework that integrates reinforcement learning with SMT (Satisfiability Modulo Theories) solving, and (3) a curated dataset of symbolic constraints. Experimental results show that WARP consistently improves performance on worst-case constraint reasoning. Leveraging the curated constraint dataset, we use reinforcement learning to fine-tune a model, WARP-1.0-3B, which significantly outperforms size-matched and even larger baselines. These results demonstrate that incremental constraint reasoning enhances LLMs’ ability to handle symbolic reasoning and highlight the potential for deeper integration between neural learning and formal methods in rigorous program analysis.
1. Introduction
Understanding the worst-case algorithmic complexity of software systems has many practical applications, ranging from compiler optimisations, finding and fixing performance bottlenecks, or improving cybersecurity in the presence of denial-of-service attacks. Profilers are typically used to find performance bottlenecks; however, they are inherently incomplete, as they are limited by the number of test inputs used. Symbolic execution techniques, such as WISE (wise-benchmarks) and SPF-WCA (spf-wca), attempt exhaustive analysis but face severe scalability limitations, due to the exponential growth of possible execution paths. Meanwhile, large language models (LLMs), exemplified by GPT-5 (openai-gpt5-systemcard), have demonstrated remarkable capabilities in a variety of programming tasks, including code generation, completion, and repair (poldrack2023ai; fu2023chatgpt). However, their sub-symbolic architecture–based on statistical pattern recognition rather than explicit logical reasoning–is not immediately suited for the kinds of rigorous analysis required in worst-case complexity reasoning or formal software verification. This gap has catalysed the field of neurosymbolic AI, which reconciles the flexibility of neural inference with the precision of symbolic reasoning. Recent work in the automated theorem proving domain has applied various approaches like reinforcement learning, retrieval-augmented generation and integration of external symbolic tools (often called “hammers”) to uplift the capabilities of LLMs in the neurosymbolic reasoning dimension (xin2024deepseekproverv15harnessingproofassistant; yang2023leandojotheoremprovingretrievalaugmented; jiang2022thorwieldinghammersintegrate).
Building on this momentum, we introduce Worst-case Asymptotic Reasoner for Path constraints (WARP), a unified neurosymbolic framework for the synthesis and verification of worst-case path constraints, as illustrated in Figure 1. WARP combines symbolic program analysis, reinforcement learning with solver-guided feedback, and rigorous benchmarking to improve the asymptotic reasoning capabilities of LLMs. Our work advances neurosymbolic reasoning for program analysis by addressing the challenge of integrating formal verification with LLMs, and it makes the following contributions:
Incremental Worst-Case Constraint Reasoning
As one of our primary contributions, we introduce an incremental inference strategy for worst-case constraint reasoning with LLMs: obtain solver-checked worst-case constraints for small input sizes using existing analysis tools, then (train and) prompt LLMs to generalise these constraint patterns to larger inputs. This reframes the task as semantics-preserving pattern induction, sharply reduces the effective search space, mitigates path explosion, and yields substantial gains over direct prompting on large inputs.
Neurosymbolic Approach for Worst-Case Complexity Analysis
At the core of the WARP framework is a novel learning paradigm, which we term Reinforcement Learning from Solver-Guided Feedback (RLSGF). This method fine-tunes LLMs to generate symbolic constraints by leveraging feedback from SMT solvers, for the synthesis of semantically valid and generalisable SMT formulas for analysing software programs. Rather than exhaustively exploring an exponentially growing search space, WARP features training on tractable, small-scale program input constraints, enabling the model to extrapolate constraint patterns to larger inputs—where traditional tools often fail due to path explosion.
Fine-Tuning and Benchmarking Datasets
To advance and evaluate worst-case–constraint synthesis, we introduce two complementary resources. First, a fine-tuning dataset of synthetically generated programs and their worst-case path constraints is designed to elevate the symbolic-reasoning skills of LLMs. Secondly, a first-of-its-kind benchmark that assesses the semantic fidelity of generated constraints, formal correctness, and generalisation. By focusing on the under-explored task of LLM-driven synthesis of worst-case constraints, the benchmark provides a rigorous, reproducible testbed for future work in neurosymbolic reasoning for WCA spectrum.
A Trained Model for Worst-Case Symbolic Constraint Reasoning
We develop and also release WARP-1.0-3B, an LLM fine-tuned using the framework to produce symbolic constraints that generalise across input sizes and outperform baselines on asymptotic prediction tasks.
2. Related Work
2.1. Symbolic Program Analysis
Early program model checkers like Java PathFinder (JPF) (jpfhavelund) systematically explore program execution states to find bugs, but face scalability issues in complex programs. Symbolic PathFinder (SPF) (10.1145/1390630.1390635) extended JPF by executing programs on symbolic inputs instead of concrete values, using constraint solvers to solve path conditions. This enabled automated test generation and analysis of Java bytecode with high coverage. Symbolic execution inherently suffers from path explosion in which the number of feasible paths grows exponentially with the number of branches. Worst-case analysis (WCA) further accentuates this problem, as it often requires the exploration of an asymptotically exponential number of execution paths, thereby rendering exhaustive analysis computationally intractable. Prior works introduced heuristics in attempt to mitigate path explosion. For example, WISE (wise-benchmarks) uses symbolic execution at small input scales and learns simple “generator” policies (e.g. always-true branches) to steer exploration toward worst-case paths at larger scales. Meanwhile, SPF-WCA (spf-wca) built as an extension to the SPF tool supports more complex branching strategies (tracking branch history) to ensure worst-case paths are explored. While these approaches prune the search, they remain limited by the exponential path growth and heavy reliance on solvers.
2.2. LLMs for Program Analysis
The emergence of LLMs pretrained on code (e.g. GPT-4, Codex) has opened new avenues for program reasoning. Modern LLMs can generate syntactically correct code and even perform simple reasoning about program behaviour in-context. This has spurred research into using LLMs for software verification and analysis tasks beyond code synthesis. For instance, (pmlr-v202-pei23a) investigated whether LLMs can infer loop invariants and other program properties important for verification. Their study demonstrated that a sufficiently powerful LLM can propose relevant invariants for synthetic Java programs, hinting at the model’s reasoning ability. Similarly, (tihanyi2024newerasoftwaresecurity) showed that LLMs can assist in vulnerability repair by generating candidate patches which are then validated with formal verification tools. Despite these promising results, purely non-symbolic approaches do not offer formal correctness guarantees. That is, an LLM’s output (e.g., an invariant or a code patch) might seem plausible while being erroneous. Consequently, because sub-symbolic LLMs operate as probabilistic next-token predictors (brown2020languagemodelsfewshotlearners), it is necessary to deploy an external checker (e.g., an SMT solver or verifier) to validate these outputs.
2.3. Neurosymbolic Reasoning
Increasingly, researchers are exploring neurosymbolic methods that integrate neural networks with symbolic reasoning to capitalise on the strengths of each. In the context of program verification, neurosymbolic systems use learning to guide or supplement traditional formal analyses. GPT-f (polu2020generativelanguagemodelingautomated) was an early landmark, integrating a transformer-based language model into the Metamath (metamath) formal system as a proof search agent. GPT-f demonstrated that training using the prover’s success to update a value function guiding search can continuously improve theorem-proving performance. Subsequent systems have expanded on this idea. For example, (zombori2021findinglongerproofs) applied deep reinforcement learning to guide a theorem prover, employing graph neural networks to represent logical formulae, with the aim of handling long, complex proofs. Techniques inspired by AlphaZero’s self-play have also been adapted to propose proof steps and receive verification feedback at each step. While effective, these RL approaches can be computationally expensive due to the need to attempt many proof paths and compute rewards from successful proofs.
An alternative direction has been to use pre-trained LLMs with post-training methods. One notable example is LeanDojo (yang2023leandojotheoremprovingretrievalaugmented), an open-source playground that enables LLMs to interact with the Lean theorem prover (lean_theorem_proving). LeanDojo provides an interface and dataset of proofs, allowing an LLM-based agent to propose proof steps and query the proof assistant. By augmenting an LLM with a retrieval mechanism over a Math library, LeanDojo can select relevant premises and tactics, significantly improving success rates in formal proofs compared to untutored LLM reasoning. DeepSeek Prover (xin2024deepseekproverv15harnessingproofassistant) pushes this synergy further by incorporating reinforcement learning signals from a proof assistant during training. DeepSeek uses Lean to provide a reward signal to fine-tune an LLM for theorem proving. In doing so, the neural model learns to align its generation with formal proof requirements, blending learned intuition with strict verification. Such efforts highlight that combining LLMs with symbolic evaluators (proof assistants, static analysers, solvers) can yield systems that are both creative and more reliable. Our work falls into this neurosymbolic spectrum, where we study and evaluate the application of LLM to synthesise worst-case complexity path constraints in SMT-LIB v2 (SMT) (BarFT-SMTLIB) format. We further formally verify that the synthesised constraints are semantically equivalent to the worst-case path constraints using the SMT solver Z3 (z3-demoura). The verification is used to guide training and evaluate model capability.
3. Problem Formulation
Symbolic worst-case analysis identifies input conditions under which software exhibits its maximal resource usage – whether time, memory, or instruction count. Identifying these conditions is crucial to prevent vulnerabilities and bottlenecks in real-world applications. However, existing symbolic execution tools, such as SPF-WCA, face significant scalability limitations due to the exponential path growth. Symbolic execution explores the program paths by solving the path constraints generated from conditional branches. The worst-case complexity emerges for input size from the paths yielding the maximal resource consumption where is the cost function for path .
3.1. Motivating Example
Consider the example QuickSort (see 1). Worst-case analysis can derive constraints for small inputs (e.g. ) succinctly represented as SMT formula:
| (assert (and (<= in0 in2) (<= in1 in2) (<= in0 in1))) |
As input size grows (e.g. ), the symbolic execution tool’s computation becomes increasingly intensive, quickly exceeding thousands of logical clauses. In particular, this represents a -fold increase in the number of execution paths compared to input above.
Even with optimised symbolic methods, like history-preserving guidance, this quickly becomes computationally infeasible at larger input sizes, requiring paradoxical amounts of computational resources and time. Analysing QuickSort at entails exploring paths, translating to impractical CPU time and memory usage, rendering exhaustive methods computationally infeasible.
| WARP Benchmark Programs | Input Size | ||||
| 2 | 4 | 8 | 16 | 30 | |
| ArrayTwister | ✗ | ✗ | ✗ | ✗ | ✗ |
| BinarySearch | ✗ | ✓ | ✓ | ✓ | ✗ |
| BinarySearchTreeHeight | ✓ | ✗ | ✓ | ✓ | ✓ |
| BinaryTreeSearch | ✗ | ✗ | ✗ | ✗ | ✗ |
| BubbleSort | ✓ | ✓ | ✓ | ✗ | ✗ |
| CaseFlipper | ✗ | ✗ | ✗ | ✗ | ✗ |
| Collatz | ✗ | ✗ | ✗ | ✗ | ✗ |
| ComplexStateMachineParser | ✗ | ✗ | ✗ | ✗ | ✗ |
| Dijkstra | ✗ | ✗ | ✗ | ✗ | ✗ |
| DizzyRamp | ✗ | ✗ | ✗ | ✗ | ✗ |
| GreedyStepper | ✗ | ✗ | ✗ | ✗ | ✗ |
| KnapsackSolver | ✗ | ✗ | ✗ | ✗ | ✗ |
| MazeSolver | ✗ | ✗ | ✗ | ✗ | ✗ |
| MergeSort | ✗ | ✗ | ✗ | ✗ | ✗ |
| NaiveFibonacci | ✗ | ✗ | ✗ | ✗ | ✗ |
| QuickSort | ✓ | ✗ | ✗ | ✗ | ✗ |
| RampUp | ✗ | ✗ | ✗ | ✗ | ✗ |
| SortedListInsert | ✗ | ✗ | ✗ | ✗ | ✗ |
| SubarraySumFinder | ✓ | ✗ | ✗ | ✓ | ✗ |
| TowerOfHanoi | ✗ | ✗ | ✗ | ✗ | ✗ |
Notes: ✓= TRUE, ✗= FALSE
3.2. Vanilla Prompting
We first explore the direct application of an LLM to generate SMT worst-case constraints for the Java programs. The LLM is given the program and it is instructed to output a syntactically correct, canonical SMT assertion (see 2). To quantify performance under vanilla prompting, we evaluated a state-of-the-art proprietary LLM, GPT-5 (openai-gpt5-systemcard), across all programs in the WARP benchmark (see subsection 4.1). For each program, we sampled five input sizes and tasked the model with generating the worst-case path constraints at each size. A generated constraint is counted as correct only if it is both: (1) syntactically valid and (2) semantically equivalent to the ground-truth worst-case constraint verified by the Z3 solver. This dual criterion ensures that correctness reflects true logical equivalence rather than superficial string similarity.
As summarised in Table 1, GPT-5 achieves only in overall accuracy. Most outputs are syntactically well-formed but fail semantic equivalence checks, underscoring the difficulty of worst-case constraint reasoning: without additional structure, even current state-of-the-art models cannot reliably generate correct constraints. These findings motivate the incremental neurosymbolic reasoning strategy of WARP: rather than requiring the model to induce full constraints from scratch, we scaffold learning with solver-verified instances at smaller input sizes, enabling systematic extrapolation to larger and otherwise infeasible regimes.
3.3. Augmenting Symbolic Execution with LLMs
We propose a neurosymbolic approach that uses SPF-WCA and LLMs in an incremental fashion. Concretely, we use SPF-WCA to obtain constraints at small input sizes and we task the LLM to:
Learn constraint patterns from those tractable input sizes.
Extrapolate these patterns to predict constraints for significantly larger inputs.
Definition 3.1 (Incremental Worst-Case Constraint Reasoning).
Given a sequence of solver-derived constraint sets for small input sizes , infer a constraint set for a larger input size such that with respect to solver-based semantic equivalence.
This incremental formulation reframes worst-case analysis as a pattern induction problem: instead of exploring all paths for large , the learner generalises from small-scale constraints to approximate or reproduce . This reduces the exponential search space, mitigates path explosion, and enables analysis at input scales where symbolic execution alone is infeasible.
4. Methodology
Figure 1 gives a high-level overview of WARP. Note that WARP includes a benchmark designed to evaluate constraint generalisation, and it is applied to finetune an LLM through reinforcement learning to synthesise SMT constraints for our neurosymbolic task. Our results show that the incremental reasoning approach significantly improves LLMs’ ability to solve worst-case constraint reasoning tasks. The fine-tuned WARP-1.0-3B model also matches the performance of much larger proprietary and open-source baselines in terms of both accuracy and semantic validity of generated constraints. Therefore, our results underscore WARP’s effectiveness as a lightweight neurosymbolic alternative.
4.1. Incremental Worst-Case Path Constraints Curation
The defining feature of WARP is its incremental approach to worst-case constraint reasoning. Instead of asking an LLM to derive large- constraints from scratch, we first curate solver-verified constraints at small input sizes where symbolic execution is efficient and reliable. These small- constraints act as building blocks: they capture the recurring structural patterns of worst-case program behaviour and provide the context needed for the LLM to extrapolate to larger, harder cases. Concretely, we use SPF-WCA (spf-wca) to analyse programs at tractable input sizes in its default configuration. SPF-WCA’s default cost model utilises the depth of the symbolic execution path, i.e., the more branches with symbolic conditions are visited during the execution, the more costly such a path is. For each program and size , SPF-WCA yields worst-case path constraints in SMT format. By collecting these constraints across multiple small- values, we build sequences that illustrate how the constraint structure evolves as increases.
WARP leverages these curated constraints in an incremental fashion during both inference and training. During inference, curated examples at smaller values are included in the prompt, and the model is tasked with generating the constraint for a larger, held-out input size. During fine-tuning, they form the foundation of a dataset that familiarises LLMs with the syntax and semantics of SMT constraints while teaching them the progression of constraint patterns across input sizes. This setup explicitly isolates the ability to generalise incrementally, ensuring that observed performance gains reflect learned symbolic reasoning patterns rather than access to ground-truth large- solutions.
4.2. Dataset Creation and Benchmarking
In order to accommodate the novel application of LLMs for the prediction of worst-case path constraints, a corpus of Java programs and the worst-case execution path constraints that follow have been collected and formatted to enable the development of generalisation and predictive capabilities in LLMs through reinforcement learning. For fine-tuning, we assembled a diverse dataset of 23,600 instances—comprising 17,700 training examples and 5,900 test examples—that span a broad range of symbolic constraints, from simple arithmetic checks to complex logical compositions. This collection includes both synthetic programs and examples derived from real-world applications. Notably, a portion of the dataset comes from SPF-WCA and the Badger tool’s examples (Noller_2018), which capture a variety of behaviours representative of real-world applications.
Independent from our fine-tuning dataset, we compiled another test set consisting of over 600 instances from 20 Java programs. Each instance includes ground-truth SMT constraints for several smaller values from a single Java program, designating a larger as the held-out target the model must synthesise (see Figure 2). By comparing the generated constraints against the held-out ground-truth, we directly measure a model’s ability to extrapolate the underlying symbolic rule rather than merely memorising observed patterns. These programs implement different algorithms, spanning array manipulation, searching, sorting, graph traversal, optimisation and recursion.
Defining Difficulty via Jumps in
We posit that difficulty depends not only on the upper bound of but also on the magnitude of the leap from the largest available candidate example to the target. For instance, employing examples at and querying can be considered more challenging than progressing through intermediate values such as () to . Larger leaps demand more extensive extrapolation, while smaller, incremental steps emphasise granular pattern replication.
Tiered Complexity and Model Limitations
To accommodate both modest and extensive leaps without incurring path explosion, we classify each instance into small, medium, and large tiers, each associated with distinct ranges of jump size. As input size increases, synthesising the corresponding worst-case constraints requires more reasoning and results in longer outputs, often nearing or exceeding the model’s maximum generation capacity. While providing more examples in the prompt can improve output quality by offering richer context, it also encourages the model to generate longer reasoning chains, which consume a greater portion of the output budget. This introduces a critical trade-off: richer prompts can improve intermediate reasoning, but increase the risk that the model truncates before fully emitting the final constraint. To mitigate this, we impose an upper limit on prompt size to ensure adequate capacity is preserved for generation. When the full set of example constraints exceeds this limit, we reduce the prompt to three representative examples—the first, the median, and the last. If the reduced prompt still exceeds the threshold, we exclude the instance entirely. Although this may conservatively filter out potentially solvable cases, it serves as a practical safeguard: if three examples already exceed the output limit, the constraint for a larger is likely even longer and prone to truncation.
Note that this constraint is not only a technical safeguard but also a deliberate design choice intended to foster broader accessibility. By bounding prompt and output sizes, we make the benchmark approachable for smaller models, enabling research on efficient reasoning under limited resources. Looking forward, we view this as a foundation for extensible benchmarks that systematically explore trade-offs between model scale, prompt complexity, and output fidelity—paving the way for evaluating both scaling trends and compact, capable alternatives.
Final Benchmark Set and Scope
After refining the benchmarking dataset through these constraints, we obtain a consolidated set of 671 instances. Each instance includes a prompt (featuring constraints at multiple smaller ) and asks for a constraint at a larger , alongside the corresponding solution. Restricting to at most 30 allows for practical verification using the Z3 SMT solver (z3-demoura), ensuring each instance can be definitively classified as correct or incorrect.
4.3. Reinforcement Learning from Solver-Guided Feedback
We also contribute WARP-1.0-3B, a 3B parameter model fine-tuned to generate SMT constraints for worst-case execution. For fine-tuning, we adapt the computationally-efficient framework proposed by sheng2024hybridflow with the following key components:
Training Template
We build upon Qwen2.5-3B (qwen2025qwen25technicalreport) and adapt it using a task-specific prompt designed to elicit SMT constraint generation. As illustrated in 3, the model is required to first reason through the problem before providing its final answer. This design choice, inspired by deepseekai2025deepseekr1incentivizingreasoningcapability, avoids imposing content-specific biases and allows the model’s intrinsic problem-solving capabilities to emerge naturally.
Reward Modelling
We adopt a rule-based reward system that consists of two types of reward:
Semantic Reward. For each predicted SMT constraint we call the Z3 solver to check logical equivalence with the ground-truth constraint . If Z3 proves , the episode receives a positive reward; otherwise it receives no reward. This binary signal is cheap to compute and correlates well with downstream verification performance described in subsection 4.4.
Syntactic Reward. A smaller bonus is granted when the output exactly matches the required template <think>…</think><answer>…</answer>. Any deviation yields zero bonus, discouraging malformed outputs without allowing the model to trade correctness for formatting.
The overall reward is a weighted sum of the two components:
This reflects our prioritisation of logical correctness over surface-level fluency. We intentionally leave the content of the <think> block unscored, focusing evaluation budget on verifiable constraint generalisation rather than natural language reasoning quality.
Reinforcement Learning Algorithm
To optimise the model efficiently, we employ Group Relative Policy Optimisation (GRPO) (shao2024deepseekmathpushinglimitsmathematical) as an alternative to standard Proximal Policy Optimisation (PPO) (schulman2017proximalpolicyoptimizationalgorithms). GRPO eliminates the need for a value model (critic) by estimating advantages from group scores, thereby reducing computational costs associated with a critic model.
Hyper-parameter Selection
For the purpose of validating our proposed approach, we adopt a default hyper-parameter settings set by vLLM (kwon2023efficientmemorymanagementlarge). Our primary objective is to demonstrate the viability of applying our framework onto LLMs, rather than to achieve optimal performance. Although more extensive hyper-parameter tuning may yield improved results, we fix the policy’s (actor) temperature, nucleus sampling (top-p) and gradient clipping at 1, as our aim is to demonstrate the feasibility of low-cost neurosymbolic training pipelines and motivate further exploration in this direction.
Training Setting
We fine-tune our pre-trained model using a constant learning rate of and a KL divergence penalty weighted at 0.001. Training is performed in bfloat16 precision on 2 NVIDIA A100 80GB GPUs with a batch size of 256 and both the input and output sequence length are capped at 2048. In our GRPO setup, we sample 16 rollouts per prompt and estimate the advantage by using the group’s reward mean and standard deviation as a baseline. These settings reflect our available hardware resources.
4.4. Evaluation and Performance Metrics
Our evaluation protocol quantifies the performance of SMT constraint generation through automated, reproducible measures:
Automated Verification
Let be the ground-truth constraint and the constraint generated by the model. We deem them correct when they are logically equivalent:
This bidirectional UNSAT test guarantees that no interpretation satisfies one constraint while violating the other, thereby certifying semantic equivalence. The resulting boolean signal is consumed by the reward function described in subsection 4.3.
Accuracy Metric
We define accuracy as the proportion of tasks for which the generated SMT constraint is both syntactically valid and semantically correct (i.e., successfully verified by Z3). By measuring solver-verified equivalence to the true worst-case execution condition, our metric eschews surrogate objectives such as BLEU or perplexity, which offer no guarantee of formal correctness.
5. Experiments
We evaluate WARP, our neurosymbolic approach for incremental worst-case constraint reasoning, across a wide range of popular large language models to assess its general applicability. In addition, we present WARP-1.0-3B, a fine-tuned model trained with the WARP approach, which achieves the best performance in its size range and even surpasses several larger open-source and proprietary baselines. Together, these evaluations demonstrate both the broad effectiveness of WARP when applied to existing models and its ability to lift performance across scales, enabling models of different sizes to reach levels typically reserved for much larger ones. The evaluation of WARP is organised around three research questions:
- RQ1.:
How does incremental worst-case constraint reasoning enable LLMs to outperform vanilla prompting, particularly when extrapolating from small input sizes to larger ones?
- RQ2.:
To what extent does WARP’s performance improvement reflect genuine symbolic reasoning ability rather than domain-specific adaptation to SMT-LIB syntax?
- RQ3.:
How robust is incremental worst-case constraint reasoning across different model scales, varying extrapolation regimes, and program families representative of real software workloads, and what practical trade-offs emerge in terms of accuracy, scalability, and computational cost?
5.1. Evaluation Setup
Evaluation Prompt
We assess each model’s ability to produce complete, structured SMT constraints by using the uniform instruction set (see 4). This instruction set directs the model to present its full reasoning between <think>…</think> tags and to return an exact, canonical SMT assertion between <answer>…</answer> tags, without any additional commentary or truncation. All evaluations are performed zero-shot, with no imposed limit on output length.
Baselines
We benchmark WARP-1.0-3B against nine open-source checkpoints (3B–32B) and several contemporaneous OpenAI models, selected to span a representative range of model scales within our computational budget.
| Model | Trial 1 (%) | Trial 2 (%) | Trial 3 (%) | Average (%) |
| Qwen2.5-3B-Instruct | 11.18 | 9.99 | 8.20 | 9.79 (±1.49) |
| DeepSeek-R1-Distill-Qwen-7B | 18.63 | 16.39 | 16.24 | 17.09 (±1.30) |
| Qwen2.5-7B-Instruct | 21.46 | 25.04 | 22.95 | 23.15 (±1.79) |
| Falcon3-10B-Instruct | 28.02 | 29.81 | 29.36 | 29.06 (±0.90) |
| GPT-4o-2024-11-20 | 37.26 | 36.96 | 33.08 | 35.77 (±2.33) |
| DeepSeek-R1-Distill-Qwen-14B | 36.51 | 36.36 | 35.32 | 36.07 (±0.66) |
| WARP-1.0-3B | 41.43 | 41.58 | 40.24 | 41.08 (±0.72) |
| GPT-4.1-2025-04-14 | 44.11 | 45.31 | 42.47 | 43.96 (±1.44) |
| Qwen2.5-14B-Instruct | 44.86 | 46.20 | 48.73 | 46.60 (±1.94) |
| DeepSeek-R1-Distill-Qwen-32B | 48.88 | 50.52 | 52.91 | 50.77 (±2.02) |
| GPT-4.1-mini-2025-04-14 | 57.82 | 59.61 | 55.89 | 57.77 (±1.86) |
| Qwen2.5-32B-Instruct | 70.19 | 70.64 | 68.70 | 69.85 (±1.05) |
| o4-mini-2025-04-16 | 86.89 | 87.78 | 90.61 | 88.43 (±1.52) |
| GPT-5-2025-08-07 | 95.83 | 95.08 | 95.50 | 95.47 (±0.38) |
5.2. Overall Results
Our overall results in Table 2 first demonstrate that incremental worst-case constraint reasoning consistently equips LLMs to outperform vanilla prompting. While GPT-5 represents the strongest proprietary baseline, its performance under vanilla prompting falls short of all models evaluated with WARP when equipped with the incremental reasoning strategy. Specifically, GPT-5 achieves only 13% accuracy when prompted directly, but reaches 95.47% under WARP, underscoring that incremental worst-case constraint reasoning is essential for enabling reliable generalisation.
The fine-tuned model WARP-1.0-3B achieves an average zero-shot generalisation accuracy of on our worst-case symbolic-constraint benchmark, evaluated over three independent trials. This performance represents a more than fourfold improvement over the instruct-tuned Qwen2.5-3B-Instruct baseline () and nearly doubles the accuracy of DeepSeek-7B (). Furthermore, WARP-1.0-3B surpasses mid-scale models such as Falcon3-10B () and DeepSeek-14B (), while maintaining a trial-to-trial variance of less than percentage point, indicating strong stability and consistent solver alignment. Table 2 presents the zero-shot generalisation results for all models, including DeepSeek-R1-Distill-Qwen variants with 7B, 14B, and 32B parameters (hereafter DeepSeek-7B, DeepSeek-14B, and DeepSeek-32B, respectively).
The recently released GPT-5 achieves an average of on our benchmark when combined with the incremental worst-case constraint reasoning strategy, whereas applied directly it performs poorly (see Table 1). In parallel, our solver-aligned reinforcement learning framework (RLSGF) produces large gains at the small scale: WARP-1.0-3B improves from (instruct variant of base model) to . These results establish both the upper bound provided by frontier models and the relative improvement enabled by RLSGF at compact scales.
LABEL:fig:model-accuracy depicts each model’s average accuracy plotted against parameter count on a log scale. Compact baselines such as Qwen2.5-3B () and DeepSeek-7B () lie at the lower end, while Qwen2.5-7B (), Falcon3-10B (), and mid-sized open-source models like DeepSeek-14B () and Qwen2.5-14B () occupy the middle. Among the largest open models, DeepSeek-32B scores and Qwen2.5-32B scores , alongside proprietary model GPT-4o () and its successor GPT-4.1 (), whose sizes are not publicly disclosed but are believed to be comparable. Notably, WARP-1.0-3B outperforms its parameter-size peers and matches or surpasses several B models. WARP-1.0’s deviation from parameter-only scaling highlights the significant impact of RLSGF, demonstrating that integrating formal SMT validation into the fine-tuning loop enables compact LLMs to synthesise semantically precise worst-case path constraints, achieving fidelity comparable to much larger models.